Suppose you pass light from a He_Ne lesser through two slits separated by a 0.01mm and you find that the third bright line on a screen is formed at an angle of 10.95 degrees relative to the incident beam. What is the wavelength of the light? What is the highest order constructive interference possible with this system?
1
Expert's answer
2020-04-24T10:51:29-0400

1)


"d\sin{\theta}=m\lambda"

"0.01\sin{10.95}=m\lambda"

"\lambda=6.3\cdot10^{-4}mm=633\ nm"

2)


"d=M\lambda"

"6.3\cdot10^{-4}M=0.01"

"M=15"